Length of the rectangle is 6m more than its width.the perimeter of the rectangle is 60m.find the length and breadtth.
Answer by lwsshak3(11628)

Length of the rectangle is 6m more than its width.the perimeter of the rectangle is 60m.find the length and breadtth
***
let x=width of rectangle
x+6=length of rectangle
2(width+length)=perimeter
2(x+x+6)=60
4x+12=60
4x=48
x=12
x+6=18
width of rectangle=12 m
length of rectangle=18 m